OpenStudy (jhannybean):

Is there a way to add more LaTeX symbols to OS? For example the "arc" symbol, and others?

TheSmartOne (thesmartone):

Umm. maybe there is?

TheSmartOne (thesmartone):

Maybe @Jaynator495 or @thomaster would know...

jaynator495 (jaynator495):

there already is a command within LaTeX that allows you to use any symbol within the Ariel Unicode MS pack... And Unicode... There Is Probably more that can even be used... SURE you may not be able to use a \symbolhere command, however you can still put it in... looks like TheSmartOne Was right to call me =P

OpenStudy (confluxepic):

How else did they make LaTeX. They added things. So there should be a way. @Jhannybean @Jaynator495 @~ @TheSmartOne

jaynator495 (jaynator495):

you cant add a command like that that doesnt exsist... but theres a command to utilize any symbol (such as that) just ell me if you want to know it xD

TheSmartOne (thesmartone):

You can use @Abhisar 's search page. It has a \(\ \LaTeX\) editor ton it.

TheSmartOne (thesmartone):

editor on**

OpenStudy (abhisar):

http://webyab.hostoi.com/search.html

jaynator495 (jaynator495):

\(\varsigma \varsigma \varpi \mu \nu \iota \varpi \mu \varpi \vartheta \vartheta \varepsilon \varepsilon \left ( \left ( \left ( \left [ \left [ \left [ \left [ \left \lfloor \left \lfloor \left \{ \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \| \left \lfloor \left \lceil \left \lfloor \sum \bigcup \oint \oint \prod_{\bigcup_{\oint_{\oint \oint \sqrt{\int_{^{^{^{^{\: \: \: \: \: \; \, \pm \pm \mp \mp \wedge \wedge \triangleleft \blacklozenge \bullet \bullet \circ \wedge \wedge \wedge \triangleright \bigcirc \bigcirc \blacklozenge \therefore \therefore \partial \jmath \Re \jmath }}}}}^{}}}^{}}^{}}^{} \right \rfloor \right \rceil \right \rfloor \right \| \right \| \right \| \right \| \right \| \right \| \right \| \right \| \right \| \right \| \right \| \right \| \right \| \L \OE \O \o \sqsubseteq \sqsubseteq \sqsubseteq \sqsupset \sqsupseteq \nsupseteq \nsupseteq \supset \supset \dot{\dot{\dot{\dot{\check{\check{\widehat{\widehat{\rightarrow \leftharpoonup \leftharpoonup \Leftarrow \Leftarrow \Leftarrow \rightharpoonup n \to x \mapsto x^2 x \mapsto x^2 }}}}}}}}\right \| \right \| \right \| \right \| \right \| \right \} \right \rfloor \right \rfloor \right ] \right ] \right ] \right ] \right ) \right ) \right )\)

jaynator495 (jaynator495):

mathjex101 using it as art pretending its latex XD

OpenStudy (anonymous):

How exactly do you use unicode in Latex?

OpenStudy (anonymous):

I guess the problem is there are symbols which are supported by latex, or in it packages, cannot be used here. To elaborate on Jhannybean's example of "arc". When we want to type "arc AB" of a circle, we can't type it. Neither the command \widearc nor \overarc works. Even you search "LaTeE arc" on the search site TheSmartOne posted, you cannot find such "editor". I guess that's why Jhannybean asked this question.

OpenStudy (anonymous):

Oh, I found in some ways to represent arc AB from e.mccormick's reply [1]: 1) \overset{\huge\frown}{AB} <-- looks better, more complicated though 2) \widehat{AB} <-- doesn't look quite right Reference: [1] http://openstudy.com/study#/updates/547b3e29e4b0749d9478388e

jaynator495 (jaynator495):

you can add any unicode symbol in latex... all you need to do is type the symbol, while your in a field that txt can be written, unfortantly openstudy doesnt accept or handle symbols very well so at first it might appear as many ?'s refresh the page and it will show correctly.